OpenStudy (pfenn1):

What am I doing wrong here? I'm trying to get all the words to NOT be italicized. \text {new number} = \frac{{\cancel {miles}\text per hour}}{\cancel{\text {miles}}}

OpenStudy (pfenn1):

\[\text {new number} = \frac{{\cancel {miles}\text per hour}}{\cancel{\text {miles}}}\]

OpenStudy (kinggeorge):

If you're wondering why the "erhour" and "miles" is italicized, you need to put it in \text brackets also. Your new thing should be \text {new number} = \frac{{\cancel {\text{miles}}\text {per hour}}}{\cancel{\text {miles}}} \[\text {new number} = \frac{{\cancel {\text{miles}}\text {per hour}}}{\cancel{\text {miles}}}\]

OpenStudy (pfenn1):

Thanks.